OpenStudy (calculusxy):

If the speedometer of a car reads a constant 40km/hr, can you say the car has a constant velocity? Why or why not?

OpenStudy (calculusxy):

@Michele_Laino This is like basic physics.

OpenStudy (alekos):

velocity is defined as speed and direction

OpenStudy (michele_laino):

I think that we can only say that the magnitude of our velocity is 40 Km/h

OpenStudy (michele_laino):

namely, the motion of our car can be a uniform circular motion

OpenStudy (calculusxy):

So is it a yes?

OpenStudy (michele_laino):

I think no, since as @alekos as well said the velocity is a vector quantity, which is defined by a speed or magnitude and a direction

OpenStudy (calculusxy):

So since it doesn't have a specific direction it cannot be considered as a "constant velocity," even though it has a constant speed?

OpenStudy (michele_laino):

yes! That's right!

OpenStudy (calculusxy):

I have more questions, since I like to make sure many things. :)

OpenStudy (michele_laino):

ok!

OpenStudy (calculusxy):

What two controls in a car can use a change in speed?

OpenStudy (calculusxy):

I wrote the gas and brake pedals. But I am not sure if I should include the steering wheel, since it might change the velocity as well.

OpenStudy (michele_laino):

I think that you have to include the steering wheel, since by means of the steering wheel we can bend the trajectory of our car, so we can change the velocity of our car

OpenStudy (calculusxy):

So do i write the two pedals (gas and brake) and the steering wheel?

OpenStudy (michele_laino):

yes! I think so!

OpenStudy (sweetburger):

I agree that the steering wheel would cause a change in direction and therefore could have affect on the velocity

OpenStudy (calculusxy):

Thanks for the continuous help. Next question: What quantity describes how quickly you change how fast you're traveling?

OpenStudy (sweetburger):

Acceleration? right?

OpenStudy (calculusxy):

Yes that's what I thought. But I am not quite sure.

OpenStudy (sweetburger):

change in velocity over time*

OpenStudy (michele_laino):

right! The steering wheel can produce a centripetal acceleration which is acting on our car @sweetburger

OpenStudy (calculusxy):

@Michele_Laino For this question:What quantity describes how quickly you change how fast you're traveling? Would the answer be "acceleration"?

OpenStudy (alekos):

Previously you asked "What two controls in a car can change the speed?" (not velocity) that would only be brake and gas pedal.

OpenStudy (michele_laino):

yes! I think so, since the acceleration, as a number, is the change in speed

OpenStudy (sweetburger):

@alekos I think you are correct considering it is only asking for change in speed.

OpenStudy (michele_laino):

you are right! @alekos

OpenStudy (calculusxy):

What is the acceleraiton of a car that travels in a straight line at a constant speed of 100km/hr?

OpenStudy (calculusxy):

the acceleration would be 100km/hr since it is travels in a straight line

OpenStudy (sweetburger):

if it is at a constant speed of 100km/hr there is no change in velocity so 0/t = 0

OpenStudy (michele_laino):

if the speed is constant, then the acceleration as a simple number, and not as a vector quantity, is zero

OpenStudy (calculusxy):

Oh okay

OpenStudy (sweetburger):

the velocity must change over a specific period of time to find a acceleration value

OpenStudy (michele_laino):

the acceleration as a vector quantity, since the motion is along a straight line, is the null vector

OpenStudy (calculusxy):

What is the acceleration of a car moving along a straight line path that increases its speed from zero to 100km/hr in 10 seconds?

OpenStudy (calculusxy):

Yes that's why i deleted it

OpenStudy (michele_laino):

we have: 100 Km/h= 27.78 m/sec so acceleration is: 27.78/10=2.778 m/sec^2

OpenStudy (calculusxy):

yeah i was overlooking the "per hour" thank you for correcting that

OpenStudy (calculusxy):

BY how much does the speed of a vehicle moving in a straight line change each second when it is accelerating at 2km/hr (two kilometers per hour per second; that is every second, the velocity is increasing two kilometers per hour.)

OpenStudy (michele_laino):

after a time t (seconds), the requested change is 2*t Km/h

OpenStudy (calculusxy):

I don't quite understand

OpenStudy (michele_laino):

by definition, acceleration a is given by the subsequent formula: \[a = \frac{{\Delta v}}{{\Delta t}}\] so we have: \[\Delta v = a\Delta t\]

OpenStudy (michele_laino):

now, a= 2(Km/h)/sec and if \Delta t= 1 seconds, then we can write: \[\Delta v = a\Delta t = 2 \times 1 = 2\frac{{Km}}{h}\]

OpenStudy (anonymous):

ansur at my fb : https://www.facebook.com/ParthKohli

OpenStudy (calculusxy):

What does \[a \Delta t\] mean

OpenStudy (calculusxy):

@Michele_Laino

OpenStudy (calculusxy):

Moving on to the next question: Why does the unit of time never twice in the unit of acceleration?

OpenStudy (calculusxy):

Would it be that the acceleration is found by dividing the velocity by time: \[(distance / time) / time => distance / time^2\]

OpenStudy (michele_laino):

\[a\Delta t\] is the change in speed

OpenStudy (calculusxy):

So how would i calculate the speed?

OpenStudy (calculusxy):

QUestion: Calculate the speed (in m/s) of a skateboarder who accelerates from rest for 3 seconds down a ramp at an acceleration of 5/m^2.

OpenStudy (michele_laino):

here we have to apply this formula: \[v = a\Delta t = 5 \times 3 = ...m/\sec \]

OpenStudy (calculusxy):

so for this question would we do: \[V_f = V_i + (acceleration \times time)\]

OpenStudy (michele_laino):

yes! nevertheless the initial speed is zero, since the skateboarder is starting from the rest

OpenStudy (calculusxy):

V_i = 0 (initial velocity) acceleration = 5 m / s^2 time = 3 seconds \[V_f = 0 + (5 \times 3)\]

OpenStudy (michele_laino):

that's right!

OpenStudy (calculusxy):

15 m/s ?

OpenStudy (michele_laino):

yes!

OpenStudy (calculusxy):

THANK YOU SO MUCH!!! Last question: Which has more acceleration when moving in straight line: car increasing its speed rom 50-60km/hr or a bicycle that goes from 0 to km/hr in the same time? Why?

OpenStudy (michele_laino):

bycicle goes from 0 to ?

OpenStudy (calculusxy):

10km/hr Sorry

OpenStudy (michele_laino):

if we consider a time interval \Delta t measured in hours, then the acceleration of the car is: \[{a_{CAR}} = \frac{{\left( {60 - 50} \right)}}{{\Delta t}} = \frac{{10}}{{\Delta t}}\;\frac{{Km}}{{{h^2}}}\] whereas the acceleration of the bycicle is: \[{a_{BYCICLE}} = \frac{{\left( {10 - 0} \right)}}{{\Delta t}} = \frac{{10}}{{\Delta t}}\;\frac{{Km}}{h}\]

OpenStudy (michele_laino):

oops.. \[{a_{BYCICLE}} = \frac{{\left( {10 - 0} \right)}}{{\Delta t}} = \frac{{10}}{{\Delta t}}\;\frac{{Km}}{{{h^2}}}\]

OpenStudy (calculusxy):

So these are both the same right?

OpenStudy (michele_laino):

sorry for my english: bicycle*

OpenStudy (michele_laino):

yes! They are the same

OpenStudy (calculusxy):

Thanks!!!!!! I honestly love your help, you have been with me for the past hour!! OMG no one, believe me does that much, except for you!

OpenStudy (michele_laino):

:)

OpenStudy (calculusxy):

Sorry, but can you explain me the answer for the question: "BY how much does the speed of a vehicle moving in a straight line change each second when it is accelerating at 2km/hr (two kilometers per hour per second; that is every second, the velocity is increasing two kilometers per hour.) "

OpenStudy (calculusxy):

I still didn't get it :(

OpenStudy (michele_laino):

the speed change as function of time interval \Delta t, is given by the subsequent formula: \[\Delta v = a\Delta t\] where a is the acceleration

OpenStudy (michele_laino):

in our case, we have: \[a = 2\frac{{Km/h}}{{\sec }}\]

OpenStudy (michele_laino):

so, if we consider a time interval \Delta t = 1 second, then after that time interval the speed change is: \[\Delta v = a\Delta t = 2 \times 1 = 2\frac{{Km}}{h}\] namely the requested speed change is: 2 Km/h

OpenStudy (calculusxy):

So the answer itself is 2km/hr ?

OpenStudy (michele_laino):

yes!

OpenStudy (calculusxy):

Thank you! And for the question with the speed from 0 ti 100km/hr in 10 seconds how did you get 27.78 m/sec

OpenStudy (michele_laino):

since we have to divide by 3.6, namely 100/3.6 = 27.78 (approximated value)

OpenStudy (calculusxy):

why 3.6?

OpenStudy (michele_laino):

more explanation, we have: \[1\frac{{Km}}{h} = \frac{{1000}}{{3600}}\frac{m}{{\sec }} = \frac{1}{{3.6}}\frac{m}{{\sec }}\]

OpenStudy (michele_laino):

therefore: \[1\frac{m}{{\sec }} = 3.6\frac{{Km}}{h}\]

OpenStudy (calculusxy):

Oh okay!

OpenStudy (calculusxy):

Thank you soo.... much! May I know what times you are available in OS so that if i need help in physics/math, I can come and ask for your help?

OpenStudy (michele_laino):

I'am available from 16:00 (Italy time zone)

OpenStudy (calculusxy):

Well I am in NY so what would be the time in EST?

OpenStudy (michele_laino):

I think 16-6= 9:00 am

OpenStudy (michele_laino):

sorry: 16-6= 10:00 am

OpenStudy (calculusxy):

oh okay so it's like 8pm in italy now?

OpenStudy (michele_laino):

yes! at the moment it is 19:57 from me

OpenStudy (calculusxy):

you use the military time a lot?

OpenStudy (calculusxy):

or the 24 hour standard clock?

OpenStudy (michele_laino):

I use the standard 24 hour clock

OpenStudy (calculusxy):

okay b/c i use am and pm. anyway, thank you so much for your help. i REALLY appreciate it :)

OpenStudy (michele_laino):

:)